# Harmony: A Joint Self-Supervised and Weakly-Supervised Framework for Learning General Purpose Visual Representations

### v1

@@ Appendix C Reproducing MaskCLIP
As of writing this paper, the code for MaskCLIP’s implementation is not yet public so we had to reproduce it ourselves. To do that, we followed the approach in their original paper [17]. Namely, we define two vision encoders, student ${E}_{V}$ and teacher $\bar{E}_{V}$, and a text encoder ${E}_{T}$. We optimize CLIP’s contrastive loss, masked self-distillation loss, and masked language modeling simultaneously.
@@ C.1 Implementation of Masked Self-distillation
$P$ is the softmax function, and we used a masking ratio of 75%, removing all mask tokens, just like in MaskCLIP [17]. For the teacher encoder, we use a momentum parameter of 0.999 that linearly increased to 0.9999 throughout training, and we use minimal augmentation with cropping of scale [0.6, 1] and normalization only. We use the same learning rate parameters in MaskCLIP, and use a loss weight of 0.05 for the masked self-distillation and MLM objectives.
@@ C.2 Ablation of MaskCLIP
Additionally, we try out different parameters and architectures to further improve the performance of the framework. Namely, we try (1) adding the iBOT head [55], which maps the ViT outputs to vectors of size 8192 using MLP layers, (2) adding the CLS-level objective from iBOT [55] (Our Equation 6), and (3) adjusting the weight for the masked self-distillation and mlm losses. We show our results in Table 7. Having no CLS objective is almost identical to the original MaskCLIP implementation, depending on if centering and sharpening from DINO [6] is done, which is not explicitly described in their paper. Removing the CLS objective and adding the iBOT head result in identical 0-shot performances, both of which are only marginal improvements. We opt to use the iBOT head because it resulted in better generalization across values for our Table 2.
@@ D.2 CLIP and MAE Augmentations
We compare different data augmentation strategies on CLIP and MAE. For CLIP, we try using the standard augmentation of randomly cropping with a scale in the range [0.4, 1] and random horizontal flipping 50% of the time. We compared this to using the global crop augmentation from DINO and iBOT [6, 55], which we described in Section A. We show the comparison in Table 9. Both strategies produce comparable accuracies for classification, but iBOT’s augmentation produces higher segmentation IoU.
@@
. Augmentation strategy INET-1K ADE20K 0-Shot FT SSEG Standard 17.0 76.6 33.8 iBOT Global Crop 16.7 76.8 35.3
@@ D.3 Random and Attentive Masking
We also tried adopting ideas from FLIP [31] and Attentive Mask CLIP [51] in an attempt to improve the efficiency of Harmony. In FLIP [31], they show that randomly masking 50% of the images being fed to CLIP’s image encoder will retain its performance, while substantially improving efficiency. Attentive Mask CLIP builds on this idea by utilizing a ViT’s attention matrix (generated using an MAE model, which is our teacher $\bar{E}_{V}$) to mask only the patches with lowest relationship to the caption description. This is calculated using the CLS token of the vision encoder, which will encode semantic information throughout training because of the nature of CLIP’s objective. Both works used significantly larger training datasets than ours, so we wanted to evaluate the ideas on a smaller scale.
@@ Appendix E Qualitative Results
Appendix F Instabilities when Training with Mixed Precision
Throughout this work, we noticed instabilities in the pre-training when using mixed precision, with the loss becoming NaN in the first three epochs of training. This issue mainly happened when combining iBOT with MAE. To solve it, we changed the epsilon value for AdamW from $1e^{-8}$ to $1e^{-6}$, and reduced the Beta 2 value from 0.999 to 0.98. This solved the issue when pre-training on CC3M. Still, we faced the same issue when pre-training a ViT-B/16 on YFCC15M. Because of that, some of our results were run in mixed precision, and others were run using full precision. To present a fair comparison, all our Harmony experiments were in mixed precision. We attach all pre-training settings, including mixed precision, in the GitHub repository.

### v3

@@ Appendix C Reproducing MaskCLIP
As of writing this paper, the code for MaskCLIP’s implementation is not yet public so we had to reproduce it ourselves. To do that, we followed the approach in their original paper (Dong et al., 2023). Namely, we define two vision encoders, student ${E}_{V}$ and teacher $\bar{E}_{V}$, and a text encoder ${E}_{T}$. We optimize CLIP’s contrastive loss, masked self-distillation loss, and masked language modeling simultaneously.
@@ C.1 Implementation of Masked Self-distillation
$P$ is the softmax function, and we used a masking ratio of 75%, removing all mask tokens, just like in MaskCLIP (Dong et al., 2023). For the teacher encoder, we use a momentum parameter of 0.999 that linearly increased to 0.9999 throughout training, and we use minimal augmentation with cropping of scale [0.6, 1] and normalization only. We use the same learning rate parameters in MaskCLIP, and use a loss weight of 0.05 for the masked self-distillation and MLM objectives.
@@ C.2 Ablation of MaskCLIP
Additionally, we try out different parameters and architectures to further improve the performance of the framework. Namely, we try (1) adding the iBOT head (Zhou et al., 2022), which maps the ViT outputs to vectors of size 8192 using MLP layers, (2) adding the CLS-level objective from iBOT (Zhou et al., 2022) (Our Equation 6), and (3) adjusting the weight for the masked self-distillation and mlm losses. We show our results in Table 9. Having no CLS objective is almost identical to the original MaskCLIP implementation, depending on if centering and sharpening from DINO (Caron et al., 2021b) is done, which is not explicitly described in their paper. Removing the CLS objective and adding the iBOT head result in identical 0-shot performances, both of which are only marginal improvements. We opt to use the iBOT head because it resulted in better generalization across values for our Table 2.
@@ D.2 CLIP and MAE Augmentations
We compare different data augmentation strategies on CLIP and MAE. For CLIP, we try using the standard augmentation of randomly cropping with a scale in the range [0.4, 1] and random horizontal flipping 50% of the time. We compared this to using the global crop augmentation from DINO and iBOT (Caron et al., 2021b; Zhou et al., 2022), which we described in Section A. We show the comparison in Table 11. Both strategies produce comparable accuracies for classification, but iBOT’s augmentation produces higher segmentation IoU.
@@
Augmentation strategy INET-1K ADE20K
0-Shot FT SSEG
Standard 17.0 76.6 33.8
iBOT Global Crop 16.7 76.8 35.3
@@ D.3 Random and Attentive Masking
We also tried adopting ideas from FLIP (Li et al., 2023) and Attentive Mask CLIP (Yang et al., 2023) in an attempt to improve the efficiency of Harmony. In FLIP (Li et al., 2023), they show that randomly masking 50% of the images being fed to CLIP’s image encoder will retain its performance, while substantially improving efficiency. Attentive Mask CLIP builds on this idea by utilizing a ViT’s attention matrix (generated using an MAE model, which is our teacher $\bar{E}_{V}$) to mask only the patches with lowest relationship to the caption description. This is calculated using the CLS token of the vision encoder, which will encode semantic information throughout training because of the nature of CLIP’s objective. Both works used significantly larger training datasets than ours, so we wanted to evaluate the ideas on a smaller scale.
@@ Appendix E Qualitative Results
Appendix F Pseudocode
We present a general pseudocode for our algorithm. More detailed implementation is shared on the GitHub repository. A backward step is performed after each loss to free the computational graph, which saves memory.
Algorithm 1 Harmony Pseudocode
1:Input:
2:${E}_{V},\bar{E}_{V}$ // student and teacher vision network
3:${E}_{T},\bar{E}_{T}$ // student and teacher text network
4:$D_{V},D_{T}$ // vision and text decoders
5:$h_{I_{d}},h_{T_{d}}$ // image and text self-distillation heads
6:$m$ // network momentum
7:$\alpha$ // hard loss weight
8:for each $batch$ in loader do
9:     $i$, $t$ = $batch$ // get image and text caption
10:     $u,v\leftarrow\mathrm{augment}(i),\,\mathrm{augment}(i)$ // global views
11:     $\hat{u},\hat{v}\leftarrow\mathrm{mask}(u),\,\mathrm{mask}(v)$ // masked image views
12:     $\hat{t}\leftarrow\mathrm{mask}(t)$ // mask text for mlm and self-distillation
13:     $u_{s},v_{s}\leftarrow{E}_{V}(u),{E}_{V}(v)$ // extract image and text embeddings using student and teacher.
14:     $u_{t},v_{t}\leftarrow\bar{E}_{V}(u),\bar{E}_{V}(v)$
15:     $t_{s},t_{t}\leftarrow{E}_{T}(t),\bar{E}_{T}(t)$
16:     $\hat{t}_{s},\hat{t}_{t}\leftarrow{E}_{T}(\hat{t}),\bar{E}_{T}(\hat{t})$
17:
18:     $\mathcal{L}\leftarrow\mathcal{L}_{D}(h_{I_{d}}(u_{s},v_{s},u_{t},v_{t}))$ // distillation loss
19:     .$\mathcal{L}.\mathrm{backward}()$ // performed at each step to save memory
20:     $\mathcal{L}\leftarrow\alpha\mathcal{L}_{\mathrm{Hard}}(u_{s},t_{s})+(1-\alpha)\mathcal{L}_{\mathrm{Soft}}(u_{s},t_{s},u_{t},t_{t})$ // soft and hard contrastive loss
21:     .$\mathcal{L}.\mathrm{backward}()$
22:     $\mathcal{L}\leftarrow\mathcal{L}_{R}(D_{V}(u_{s},v_{s}),i)$ // pixel reconstruction loss
23:     .$\mathcal{L}.\mathrm{backward}()$
24:     $\mathcal{L}\leftarrow\mathcal{L}_{M}(D_{T}(\hat{t}_{s}),t)$ // masked language modeling loss
25:     .$\mathcal{L}.\mathrm{backward}()$
26:     $\mathcal{L}\leftarrow\mathcal{L}_{TD}(h_{T_{d}}(\hat{t}_{s}),h_{T_{d}}(t_{t}))$ // text-distillation loss
27:     .$\mathcal{L}.\mathrm{backward}()$
28:
29:     update(${E}_{V}$) // backpropagation based on the above losses
30:     update(${E}_{T}$)
31:     $\bar{E}_{V}.\mathrm{params}\leftarrow m\,\bar{E}_{V}.\mathrm{params}+(1-m),{E}_{V}.\mathrm{params}$ // EMA teacher update
32:     $\bar{E}_{T}.\mathrm{params}\leftarrow m\,\bar{E}_{T}.\mathrm{params}+(1-m),{E}_{T}.\mathrm{params}$
33:end for
Appendix G Comparison with DINOv2
DINOv2 (Oquab et al., 2024) is an obvious comparison with Harmony, as it is an improved version of iBOT (Zhou et al., 2022). However, we do not include it in our results, because we observed poor performance when pre-training DINOv2 on CC3M (linear probing accuracy below 25%), using the publicly shared codebase. Given that the modifications introduced by DINOv2 beyond iBOT are relatively minor, with many being just hyperparameter changes (see Table 1 in Oquab et al. (2024)), we would expect similar performance. We suspect a bug may be present in the code, though we were unable to identify it. As a result, we felt it would be more fair to exclude this comparison.